Find the slope of the line containing each given pair of points. If the slope is undefined, state this.
(0,9) and (-5,0)



Answer by dabanfield(803)

Find the slope of the line containing each given pair of points. If the slope is undefined, state this.
(0,9) and (-5,0)

The slope of the line, m, is the difference in the y values divided by the difference in the x values so:
m = (0-9)/(-5-0) = -9/-5 = 9/5